OpenStudy (lena772):

Calculate the theoretical mass (in grams) of Sodium Sulfate produced using only the volume and molarity of Sulfuric Acid. Calculate the theoretical mass (in grams) of Sodium Sulfate produced using only the mass of Sodium Carbonate. What is the theoretical yield (in grams) of Sodium Sulfate? What mass (in grams) of the excess reactant is left over after the reaction?

OpenStudy (lena772):

\[H_{2}SO_{4}+Na_{2}CO_{3} \rightarrow H_{2}CO_{3} + Na_{2}SO_{4}\]

OpenStudy (anonymous):

Alright this is molarity question

OpenStudy (anonymous):

First off this is a double displacement question and I assume you know it.

OpenStudy (lena772):

M= moles solute/L of solution moles solute= M*L m=4.9M*0.00698L =0.034202*105.99 g/ 1 mol = 3.6 mol However, I get that this is incorrect

OpenStudy (lena772):

That should say 3.6g not mol.

OpenStudy (anonymous):

This is a very simple question. You ask your periodic table "hey what's up man tell me what your molar mass label has to say and I'll add them up happy to the kingdom come"

OpenStudy (anonymous):

Oh shiz I didn't note "using only the mass for the reactant" ok

OpenStudy (anonymous):

You do the happy stoichometry

OpenStudy (lena772):

Do you see where I went wrong?

OpenStudy (anonymous):

Why don't you just use the stoichiometry ratio?

OpenStudy (anonymous):

I recommend stoichoimetry for the sake of simplicity because it's easier

OpenStudy (anonymous):

@Nnesha what's your take on this?

OpenStudy (lena772):

I don't know what you mean by the stoichiometry ratio.

OpenStudy (anonymous):

I am gonna give you the kingdom come so the ratio between sulfuric acid used and sodium sulfate used is 1:1 meaning that every mole of sulfuric acid results in one mole of sodium sulfate. So find the overall gram of sodium sulfate by looking at periodic table and add them together.

OpenStudy (anonymous):

for 1 mole of reactant results in 1 mole of product.

OpenStudy (anonymous):

pretty simple isn't it?

OpenStudy (anonymous):

@empty if you could elaborate what I mean that's great

OpenStudy (anonymous):

that dude got scared off

OpenStudy (anonymous):

ok I am here with you no worries

OpenStudy (anonymous):

but I am really sleepy right now vote on my thread

OpenStudy (lena772):

Add what together? I don't understand.

OpenStudy (anonymous):

@empty pitch in dude don't be the spectator ion